Meeting Summary Notes

From Cuawiki
Jump to: navigation, search

NPS COLLABORATION MEETING - FEBRUARY 1-2, 2021

HALL C UPDATE (Steve)

  • HMS status
  • many detectors have been refurbished
  • dipole momentum range up to 7.4 GeV/c (high-end momentum range of HMS) - may need to rope off platform at high fields
  • First opportunity to run NPS Phase-1 in 2023 (E12-13-010/E12-13-007 and E12-06-114 additional DVCS days moved from Hall A)
  • Estimated time to install NPS start to finish: 3-4 month assuming pre-installation tasks have been done ahead of time
  • HB removal ~ 2 weeks
  • Target foes not have to be removed to install the NPS (see page 4 of Paulo's talk
  • Action item:
  • check on plans for testing HMS hodoscope paddle in magnetic field
  • detailed installation plan
  • check what is needed formally to add E12-06-114 to the run plan, e.g., amendment to ERR 2019, submission of beam time request, etc.


MECHANICAL DESIGN (Paulo)

  • Much progress with design and procurements
  • Biggest challenge at the moment: routing cables to patch panel and accommodating movement for configuration changes
  • detector slide cart weight ~3000 lbs - need to determine how to best push it for different detector locations
  • Stair case: for any detector support structure that people will stand on need approval from structural engineer - e.g. QWeak platforms in past
  • Action items:
  • make small table with time estimates for rotating magnet, moving magnet from one side to another, etc.
  • Plan any work on target platform before May 2022 (before nuclear target running)


SENSORS/HV BOARDS&CABLES/SLOW CONTROLS (Aaron)

  • Much progress - 1100 HV cables were fabricated
  • HV cable testing still needs to be done
  • Certification of Samtec connectors
  • connector rating from manufacturer seems to not match the operational condition with range up to 1100 V
  • note: if change connector also need to change HV board
  • path forward: test connector up to 1100 V and document the results (DSG)
  • Long cables (140' GPW 42cond) will be available over time in the test lab - length selected to be what is needed in Hall C
  • Action items:
  • test HV cables
  • test Samtec connectors and document


DAQ/ELECTRONICS (Brad S.)

  • Support hardware
  • Cable run - working with Paulo
  • Grounding - isolate NPS chassis so that do not contaminate the system
  • Reminder of trigger: generated by VTP and masked to select crystals that are part of the cluster
  • Cluster trigger 3x3 with row at the seam - going to smaller cluster is possible, but involves significant work
  • VTP logic - if find a valid cluster then form trigger
  • full waveform for all crystals participating in the cluster (identified by the VTP) will be recorded
  • DAQ hardware and firmware/SW
  • FADCs are available, but need to coordinate with other projects (e.g. SBS) --> reserve FADCs for NPS
  • FADC synching with RF - never been done for FADC, might be an interesting run plan item, but not needed for NPS
  • Firmware development ongoing
  • Need Analyzer support for event-blocking mode, decoder updates - critical path item
  • examples may be available from Compton polarimeter activities
  • Slow Controls
  • LED board design nearly done - next need to procure and fabricate
  • LED control GUI needs to be designed
  • NPS Assembly area
  • Detector assembly area identified and cleared in the test lab - accross from NPS crystal clean room; some coordination with design and/or Walter's groups
  • DAQ in the test lab should be available later in the year
  • Action items:
  • Document electronics/DAQ items in the Wiki - material that could be of interest in future reviews
  • Prepare an assembly plan for CH similar to what was done for mechanical mounting


FEG UPDATES (Chris Cuevas)

  • Need to get started on sparsification firmware/SW
  • Slow controls
  • LED board - should have pulse mode available during regular run
  • VME crate should be available for LED drivers
  • Action items:
  • Start on sparsification firmware - critical path item
  • continue all other activities


ANODE CURRENT (Fernando)

  • Anode current setup is now ready - next need to develop the procedure to measure anode current
  • need personnel from NPS collaboration (2-4 hours/day)
  • Action item:
  • measure and document anode current - directly related to availability of bases - critical path item


DETECTOR FRAME (Emmanuel)

  • Assembly completed in Orsay and frame was shipped to and received at JLab in January
  • Detailed NPS mounting plan document was developed
  • Note: the final frame for the NPS will be available for the assembly in the test lab
  • Add cooling/pressure tests early on, e.g. near day 2
  • note: if disassemble and then reassembly tubes have to do a leak check everytime
  • May want to add explicitly to check for bubbles in optical grease
  • Check of magnetic shielding to be done later - need full field
  • Time frame assembly - current estimate: later in the year
  • International travel likely not possible until September
  • Need availability of and timeframe for components
  • Note: NPS assembly moves to Hall C as one unit
  • Action items:
  • Add cooling/pressure tests, etc. in NPS mounting plan
  • Check Kapton tape - isolation configuration


CRYSTAL STACKING (Vladimir, Hamlet, Hakob)

  • Estimate to have 1164 Crytur crystals at JLab in August 2021
  • some Crytur crystals may be needed for ongoing R&D projects
  • there are also on order 400 SICCAS crystals - there are less radiation hard and not as uniform as Crytur crystals
  • Crystal stacking strategy: 3 possible options based on October 2020 discussions)


MAGNET MAPPING (Charles)

  • Measurements were done up to 250A - see differences betweem Opera calculations vs. data
  • Note: measurements were done with main coil only, while Opera assumed both coils (main and corrector)
  • could rerun TOSCA model to give confidence that the magnet works
  • Magnet steel
  • magnet steel predominantly 1010
  • steel used: from BNL and additional steel from another JLab magnet
  • in general, newer steel may saturate earlier than older steel, but does not seem an issue here
  • Overall plan for the magnet:
  • Demonstrate it works as advertised
  • Fringe fields - to be tested in the hall - need accelerator test plan (HMS optics, beam displacement, etc.)
  • Action items:
  • Rerun Opera with just the main coil at 250A
  • Accelerator test plan (Jay)


SOFTWARE DEVELOPMENT (Carlos Y., Steve)

  • Working in NPS clustering algorithm - integration into Hall C Analyzer
  • idea is to develop multiple different algorithms to compare
  • Other items to be done:
  • VTP decoder updates
  • solve and implement data unblocking
  • merge multi-threaded podd with hcana
  • Data unblocking
  • There are actual data files (Brad S.)
  • Some codes are around, but nothing yet of production quality
  • some examples from Compton polarimeter activities from about one year ago are available (Alexandre)
  • Waveform analysis
  • Orsay has some examples of code
  • Data are available from SHMS
  • Calibration and monitoring software
  • need guidance on how calibrations are done - calibration methods
  • Action items:
  • Algorithm development
  • Multiblock


ASSEMBLY PLAN - ITEMS NEEDED

  • Installation plan
  • Mechanical mounting plan
  • Crystal wrapping plan
  • Software test stand and counting house assembly plan
  • Cabling maps
  • Resources plan - on-site personnel
  • Accelerator test plan


CALIBRATION PLAN

  • Calorimeter calibration Methods:
  • Cosmics
  • Elastics
  • Exclusive pi0
  • LED pulses (LED peak wavelength 465 nm)
  • Tests before moving to hall - in general test beforehand as much as possible to identify possible bottlenecks:
  • Mu metal shield test - could use Helmholtz coil setup Hall A
  • HV calibration PMT
  • DAQ/electronics - LED, cluster trigger checking
  • Cooling tests
  • Slow controls, interlocks
  • Tests after moving to hall:
  • Mu metal full assembly with magnet on
  • Fringe fields
  • Calibrations crystal
DAQ/electronics - OR w/ VTP


WACS - NPS PHASE-2

  • There are two WACS experiments: 1) unpolarized (cross section), 2) Polarized WACS (asymmetry, needs also CPS)
  • Unpolarized WACS was part of the ERR 2019 - main difference to Phase-1 is the radiator
  • For polarized WACS need additional crate for GEM readout
  • Action items:
  • prepare beam time request for unpolarized WACS
  • set milestones for formal ERR request for polarized WACS